### v2

# GIVT: Generative Infinite-Vocabulary Transformers

Abstract
We introduce generative infinite-vocabulary transformers (GIVT) which generate vector sequences with real-valued entries, instead of discrete tokens from a finite vocabulary. To this end, we propose two surprisingly simple modifications to decoder-only transformers: 1) at the input, we replace the finite-vocabulary lookup table with a linear projection of the input vectors; and 2) at the output, we replace the logits prediction (usually mapped to a categorical distribution) with the parameters of a multivariate Gaussian mixture model. Inspired by the image-generation paradigm of VQ-GAN and MaskGIT, where transformers are used to model the discrete latent sequences of a VQ-VAE, we use GIVT to model the unquantized real-valued latent sequences of a VAE. When applying GIVT to class-conditional image generation with iterative masked modeling, we show competitive results with MaskGIT, while our approach outperforms both VQ-GAN and MaskGIT when using it for causal modeling. Finally, we obtain competitive results outside of image generation when applying our approach to panoptic segmentation and depth estimation with a VAE-based variant of the UViM $*{}^{*}$ork done as a Student Researcher at Google DeepMind.
${}^{\circ}$ignificant technical contributions.
1 Introduction
After becoming the dominant architecture in natural language processing shortly after their introduction, Transformers [67] have also recently become very popular in computer vision [16, 57, 37]. Dosovitskiy et al. [16] showed that by splitting images into sequences of patches, linearly embedding those patches, and then feeding the resulting sequence of features to a transformer encoder leads to powerful image classifiers that outperform CNN-based architectures at large model and data scale. This strategy is now standard for many discriminative vision task including classification [16], detection [37], and segmentation [57].
Figure 1: We compare the standard discrete-token generative transformer (left) to our continuous, infinite-vocabulary variant (GIVT, right), using the same decoder-only architecture. At the input, GIVT linearly embeds a sequence of real-valued vectors instead of discrete tokens via lookup. At the output, GIVT predicts the parameters of a continuous distribution over real-valued vectors rather than a categorical distribution over a finite vocabulary.
Figure 2: Selected samples from the GIVT-Causal for 10 ImageNet classes (130, 130, 138, 144, 933, 145, 360, 207, 829, 248).
However, mapping an image to a sequence of (unquantized) feature vectors is not suitable for transformer-based image generation since generative transformers decoders consume and predict discrete tokens from some fixed, finite vocabulary. Such a structure naturally fits natural language, for which decoder-only models enable powerful sequential generative modeling and efficient training via teacher forcing [67, 47]. To harness these capabilities for images, recent works [18, 6, 7, 36, 43] have employed a two-stage approach which first learns a Vector-Quantized Variational Autoencoder (VQ-VAE) [66] to map images to a sequence of discrete tokens, and then learns a transformer decoder to model the latent discrete-token distribution. Another advantage of such VQ-VAE-based image tokenization is that it enables interleaved multimodal generative models, simply by concatenating the vocabularies of the different modalities including text and images [1, 26, 2].
While this two-stage approach was successfully applied to image and multimodal generation, it has several issues. The vocabulary size in VQ-VAE determines how much information can be stored in the latent code sequence, and how much of the visual modeling is done by the VQ-VAE decoder. Having a small vocabulary can make the latent modeling easier but also makes the latent code less informative, which prevents control of the low-level details in image generation, and impacts quality when using the tokens for dense prediction [30, 39] or low-level discriminative tasks [1, 26]. To mitigate this issue, the vocabulary size can be increased, which, however, can lead to low vocabulary utilization [43] so that high-fidelity VQ-VAE setups rely on a range of advanced techniques, such as entropy losses [6] or codebook-splitting [30]. Furthermore, large vocabularies lead to correspondingly large embedding matrices consuming a lot of memory, which can be an issue particularly in multimodal contexts with vocabularies combined across modalities.
To avoid these issues, we propose a modification to decoder-only transformers which removes the need for discrete tokens and therefore fixed, finite vocabularies. Specifically, we propose—to our knowledge for the first time—a generative transformer decoder which operates on sequences of real-valued vectors. As real-valued vectors can be seen as an infinite vocabulary, we call this a Generative Infinite-Vocabulary Transformer (GIVT). We make two minimal changes compared to the standard transformer decoder architecture [67, 47], see Fig. 1: 1) at the input, rather than using a sequence of discrete tokens to look up a finite vocabulary of embeddings, we linearly embed a sequence of real-valued vectors; and 2) at the output, rather than predicting the parameters of a categorical distribution over a finite vocabulary (via logits), we predict the parameters of a continuous distribution over real-valued vectors. We train this model in the same way as standard transformer decoders: with a causal attention mask and teacher forcing [67], and alternatively also explore fast progressive masked-bidirectional-modelling as in MaskGIT [12, 6, 7].
While GIVT can, in principle, be applied to any sequence of feature vectors, it can be challenging to directly model sequences that are prohibitively long or follow a complex distribution (e.g., the sequence of RGB pixels obtained by flattening a high-resolution image). Therefore, similar to the two-stage approach with VQ-VAEs and analogous the two-stage approach of latent-diffusion models [50, 46], we first learn a lower-dimensional latent space with a Gaussian-prior VAE, and then model it with GIVT.
Additionally, we port various inference schemes from the sequence-modeling literature (e.g., temperature sampling, classifier-free guidance). Surprisingly, this leads to a model which is comparable to or better than VQ-based approaches, by purely relying on real-valued tokens.
Our main contributions can be summarized as follows:
  1. 1.
    We show that GIVT matches or outperforms the standard discrete-token transformer decoder on both image generation and dense prediction tasks like semantic segmentation and monocular depth estimation (via UViM [30]).
  2. 2.
    We derive variants of standard sampling techniques for the continuous case, such as temperature sampling, beam search, and classifier-free guidance (CFG, [23]), and demonstrate their effectiveness.
  3. 3.
    We study the relationship between the degree of VAE latent-space regularization (via KL-term weighting [27, 22]) and the resulting properties of GIVT.
We emphasize that the VAE and GIVT training only relies on standard techniques from the deep-learning toolbox, and not the advanced training techniques of the VQ-VAE literature like auxiliary losses [66, 6] on the latent representation, codebook reinitialization [34], or dedicated optimization algorithms [30, 25].
Figure 3: GIVT-Causal training and inference. Left: During training, we sample a sequence of real-valued latent vectors from the VAE encoder, and train GIVT via teacher forcing. Right: During inference, we sample a sequence of vectors (left-to-right) and feed it to the VAE decoder. We note that we also explore MaskGIT-like GIVT models not shown here. No component uses a quantizer.
2 Related work
VQ-VAE for visual data tokenization  Following the success of pixel-space autoregressive modeling [65, 53, 40, 8] for image generation, moving the autorgressive modeling to the latent space of VQ-VAEs [66, 49] emerged as a more efficient alternative. The use of GANs and perceptual losses for VQ-VAE training as well as modern causal [18, 72, 68] and masked [6, 7, 36] transformers for latent modeling led to substantial quality improvements. Another active area leveraging VQ-VAEs is interleaved multimodal generative modeling of images and text [1, 26, 2]. Further, VQ-VAEs are a popular choice to tokenize the label space of dense prediction vision tasks [30, 39]. Finally, some language-inspired techniques for self-supervised learning from images rely on VQ-VAE representations [3, 70, 36].
Discretized mixtures of distributions  replace the dense prediction of the logits of a categorical distribution with a continuous mixture model which is subsequently discretized. This approach was proposed in [53] for pixel-space autoregressive modeling, to reduce the number of model parameters and to improve learning efficiency, and is also popular in neural compression [41, 10, 42].
Continuous outputs in NLP  A popular approach to handle large vocabularies in machine translation is to predict language tokens via their word embeddings with a continuous distribution, instead of token IDs with a categorical distribution [31, 32, 58, 59, 35]. Decoding is usually done in greedy fashion with embedding lookup and hence does not produce diverse samples. Further, the models consume and predict word embeddings form a fixed, finite set.
VAEs with learned priors  A rich body of literature studies improving VAEs with learned priors: Inverse autoregressive flows emerged as a popular choice [28, 9]. Other approaches are based on normalizing flows [63] or a mixture of variational posteriors with pseudo-inputs [60]. For VAEs with discrete (non-VQ) latents, learned priors based on Restricted Boltzmann Machines were studied [64, 51].
Time-series modeling with Transformers  A variety of works has recently explored transformers for time-series modeling/forecasting. Those works either use a regression loss [74, 45, 33, 20, 11], quantile forecasting [17, 38], or resort to discretizing/binning the data [48]. Somewhat related, [44, 69] regress continuous speech features from discrete tokens. None of these models predict a continuous distribution like GIVT that allows for autoregressive generation.
3 Generative infinite-vocabulary transformers
As mentioned in Sec. 1, our method is conceptually similar to recent works that train decoder-only transformer models on the discrete codes of VQ-VAEs [18, 71, 6, 7], with the crucial difference being that we do not quantize (i.e., do not use VQ). We now describe the components of our method.
3.1 VAE training
We first train a continuous $\beta$-VAE [22] with Gaussian encoder and prior as originally proposed by [27]. Given an input image $x$, the encoder $E$ predicts mean $\mu$, and variance $\sigma$ of a multivariate normal distribution with diagonal covariance matrix, and samples a representation $z$ from $\mathcal{N}(\mu,\sigma)$ using the reparameterization trick [27]. The VAE decoder then maps the latent sequence back to an image. Since we use a Gaussian encoder distribution, the KL-term in the evidence lower bound (ELBO) [27] can be computed in closed form as described in [27, Sec. F.1]. As for the reconstruction/likelihood term in the ELBO, we rely on a mixture of MSE, perceptual loss and GAN loss for image generation following [6, 18], or the categorical cross-entropy for dense prediction tasks [30]. Our encoder spatially-downsamples $x$, whereby we obtain a $z$ with spatial dimensions $h\times w$ and feature dimension $d$, with $h{=}\lceil H/16\rceil,w{=}\lceil W/16\rceil$, given a $H{\times}W$ input $x$.
The hyperparameter $\beta$ multiplying the KL-term controls how strongly $z$ is regularized. As we shall see in Sec. 5, this regularization of the VAE is important to be able to model the resulting (true) latent distribution $p(z)$ well.
[…]
GANs BigGAN-deep [5] $\numprint{6.95}$ $\numprint{0.87}$ $\numprint{0.28}$
StyleGAN-XL [54] $\numprint{2.30}$ $\numprint{0.78}$ $\numprint{0.53}$
Diffusion ADM [13] 250 $\numprint{10.94}$ $\numprint{0.69}$ $\numprint{0.63}$
Models ADM-G [13] CG $=1.0$ 250 $\numprint{4.59}$ $\numprint{0.82}$ $\numprint{0.52}$
DiT-XL/2 [46] 250 $\numprint{9.62}$ $\numprint{0.67}$ $\numprint{0.67}$
DiT-XL/2-G [46] CFG $=1.50$ 250 $\numprint{2.27}$ $\numprint{0.83}$ $\numprint{0.57}$
Sequence VQGAN [18] Top-k $=$ Mixed ${}^{\star}$ $\numprint{17.04}$
Models VQGAN [18] Top-k $=600$, CG $=0.05$ ${}^{\star}$ $\numprint{5.20}$
[…]
Table 1: Results on class-conditional $256{\times}256$ ImageNet, where GIVT-Causal models outperform their quantization-based counterparts. We report FID as well as precision and recall (where available). We use the standard ADM evaluation suite, where FID is calculated w.r.t. the training set. CG: Classifier guidance acceptance rate or scale, CFG $=w$: Classifier-free guidance with weight $w$ [23], DB-CFG $=w$: Our distribution based CFG variant (Sec. 3.3.2), Top-k: Top-k sampling [19] (“mixed” refers to multiple $k$), Beam $=B/F$: Beam search with $B$ beams, $F$ fans, $t$: Temperature sampling by scaling the predicted $\sigma$ of our models with $t$, $t_{C}$: Choice temperature for MaskGIT. Steps shows inference steps. Additional comments: ${}^{\dagger}$umbers obtained by us from public code, ${}^{\star}$nference uses activation caching.
3.2 GIVT training
We next train a GIVT to predict $p(z)$ or $p(z|c)$ (when a conditioning signal $c$ is available, e.g., in class-conditional generation). We flatten $z$ into a $hw$-length sequence of $d$-dimensional real-valued vectors (or “soft tokens”). Note how this differs from the standard VQ-VAE-based setup, where the latent transformer decoder models a $hw$-length sequence of integers denoting codebook indices. Apart from this change, we only perform two small changes to the standard transformer decoder-only architecture (see Fig. 1): We replace the embedding lookup tables at the input with a single linear layer to project from $d$ to the transformer hidden dimension. At the output, we do not predict a categorical distribution, and instead let the transformer predict the parameters of a continuous distribution.
Assuming channel-wise independence, we can model each of the $d$ entries with a $k$-mixture GMM. The GIVT model then predicts $3dk$ parameters ($dk$ means, scales, and mixture probabilities) that we interpret as $d$ scalar GMMs. Experimentally, we found it beneficial to normalize the mixture probabilities with a softmax activation, and the variance parameters with softplus.
We use the standard cross-entropy loss (which is equivalent to the negative log-likelihood) on the distribution $\tilde{p}$ predicted by GIVT, and minimize $\mathcal{L}_{\text{T}}=\sum_{c}\mathbb{E}_{z}\left[-\log\tilde{p}(z|c)\right]$, assuming the the classes or conditioning signal $c$ uniformly distributed (see Appendix A.1 for details on the loss). We train two types of GIVT models, as described next.
GIVT-Causal  Here, $z$ is flattened into a sequence line-by-line, and the GIVT predicts every $d$-dimensional vector in the sequence conditioned on all previous vectors, masking the self-attention layers to be causal [18, 67]. This training strategy is also called teacher forcing and is analogous to the latent modeling in VQ-GAN [18]. For class-conditional image generation we prepend a [CLS] vector to the input sequence, i.e., a learned vector for each class $c$.
GIVT-MaskGIT  Here, as in MaskGIT [6], we mask the input randomly during training and then gradually uncover the masked tokens during inference. The only changes compared to [6] are related to our real-valued tokens: since we have infinitely many tokens, there is no obvious choice to define a special mask token (when using VQ, one can just extend the vocabulary to contain special tokens, such as [MASK]). Instead, given $z$ and a mask $M$ indicating for every spatial location whether it is masked, we first replace the locations in $z$ corresponding to $M$ with zeros (to remove information), and then embed it with a single dense layer, as above. Additionally, we concatenate one of two learned special vectors in the feature dimension, a [MASK] vector for masked locations, and a [UNMASK] vector otherwise (we half the dimension of the embedded inputs and special tokens s.t. the final hidden dimension remains unchanged).
3.3 Inference
Given a VAE and GIVT trained as above, we run GIVT during inference either causally (see Fig. 3) or as in MaskGIT [6]. We now investigate the various inference schemes that have been developed for discrete transformers, and derive their continuous counterparts.
3.3.1 Temperature Sampling, Nucleus Sampling,
Beam Search
In causal sequence models for text (see [24] for an overview and discussion) and VQ-GAN-based approaches, it is common to tune the sampling algorithm. We start with temperature sampling, which for discrete models adapts the softmax temperature of the categorical distributions predicted at each decoding step. For GIVT, we instead scale the standard deviation of the predicted Gaussian distribution(s) and call this strategy “variance scaling”. As we will see in Sec. 4, this simple change can have a significant impact on sample quality.
Nucleus sampling [24] proposes to collect the largest logits such that its cumulative probability after normalization exceeds a threshold (for example 0.8), and to sample from this reduced-support distribution. In GIVT, for a single Gaussian prediction per output dimension, this can be approximated by truncating the predicted distributions per dimension (thereby choosing a higher-density support). We observed that this has a similar effect to variance scaling and therefore do not pursue this strategy.
We also consider beam search, which is the same for GIVT as it is for discrete transformer decoders. For every sample, we maintain $B$ beams, and at every step we sample a number of candidates for every beam (we call these “fans” here). We then compute the cumulative log probability for all beams and fans up to the current sampling step, and select the $B$ beams with the highest cumulative log probability. Finally, there is no analogous concept for top-k sampling [19] in GIVT, because it predicts continuous distributions.
Figure 4: VAE ablation: We vary the KL weight $\beta$ and number of channels $d$ when training the VAE. Dashed lines show the sampling FIDs obtained with a Base-size GIVT-Causal. As we increase $\beta$, the sampling FID improves, but the reconstruction FID increases, lower bounding the best possible sampling FID. For the main results we use $\beta=0.00005$ and $d=16$.
3.3.2 Distribution-Based Classifier Free Guidance
In the diffusion literature, classifier-free guidance (CFG) [23] has been employed with great success. Concretely, conditional diffusion models are trained with an additional null class $\emptyset$ to learn the unconditional data distribution. Then, during inference, the conditional log density is “moved away” from the unconditional one: given a guidance weight $w$, the updated (diffusion) score estimate is is obtained as
$\tilde{\epsilon}(z,c)=(1+w)\epsilon(z,c)-w\epsilon(z,\emptyset),$ (1)
where $\epsilon$ estimates the gradient of the log density of the data distribution, $\epsilon(z,c)\propto\nabla_{z}\log p(z|c)$ (see [23, Sec. 2]). From this, we now derive a CFG variant for our GIVT, since we directly predict a density. We term this approach “Density-Based CFG” (DB-CFG). Eq. 1 can be written as
$\tilde{\epsilon}(z,c)$ $\propto(1+w)\nabla_{z}\log p(z|c)-w\nabla_{z}\log p(z|\emptyset)$
$\propto\nabla_{z}\log\left(p(z|c)^{1+w}p(z|\emptyset)^{-w}\right),$
i.e., $\tilde{\epsilon}$ estimates the log of the density $p_{\text{CFG}}(z|c)\propto p(z|c)^{1+w}p(z|\emptyset)^{-w}$ (see Fig. 5). Thus, we want to adapt our models to sample from $p_{\text{CFG}}$. To this end, we follow [23] and train GIVT with an additional null class $\emptyset$. During inference, we run GIVT twice at every step, once conditional on the actual label $c$ and once conditional on $\emptyset$. To implement classifier-free guidance, we then have to sample from an unnormalized version of $p_{\text{CFG}}(z)$ derived from the two GIVT predictions. To this end, we turn to rejection sampling, which requires: 1) an unnormalized density; 2) a good proposal distribution $p^{\prime}$, that is close to the true target distribution; and 3) a scaling factor $K$ to bound the likelihood between $p^{\prime}$ and the unnormalized target density.
Figure 5: Visualization of our Density-Based Classifier-Free Guidance (DB-CFG). We show the conditional and unconditional PDFs predicted by GIVT, and the resulting CFG PDF for two values of $w$. Note how the CFG distributions become more peaked. We use rejection sampling to sample from $p_{\text{CFG}}$ (see Sec. 3.3.1).
VAE Sample
Step 1
Step 4
Step 8
Step 16
Figure 6: First Column: Two images obtained when sampling from the VAE prior, resulting in a soup of low-level image textures. Remaining Columns: Visualizing the output of GIVT-MaskGIT, for two ImageNet classes (947, 94), after 1, 4, 8, 16 inference steps. As expected, the samples start to become more coherent as we perform more inference steps.
In our setup, the distributions we mix are all Gaussian. We find empirically that the unconditional distributions (i.e., distributions predicted using the $\emptyset$ label) tend to have larger variance than the conditional ones (as visualized in Fig. 5). It is thus sensible to pick sample proposals from $\mathcal{N}(\mu_{c},2\sigma_{c})$, where $\mu_{c},\sigma_{c}$ are the parameters predicted by GIVT when given the label $c$. We empirically find that drawing 1000 samples is enough to find at least one valid sample 99.9% of the time. For the remaining ${<}0.1$%, fall back to sampling from $\mathcal{N}(\mu_{c},\sigma_{c})$.
We emphasize that the overhead of DB-CFG is small: it requires two forward passes (per inference step) instead of one to predict the conditional and unconditional distribution. We then draw 1000 samples from those in parallel on an accelerator, which is very fast. We refer to App. 3.3.2 for Python code.
4 Experiments
We explore various tasks to show the generality of GIVT.
4.1 Image generation
We use ImageNet and explore class-conditional generation (where we condition our GIVT on class labels) for 256px and 512px, as well as unconditional generation for 256px.
VAE  We closely follow the setup of MaskGIT [6]. We use their VAE architecture, built of ResBlocks (as detailed in App. A; encoder and decoder have a combined 53.5M parameters), remove the VQ layer and related losses, and replace it with a linear layer predicting $\mu,\sigma$ (Sec. 3.1). We use the same weights for reconstruction, perceptual, and GAN-loss, as well as identical optimizer parameters, as in [6, 43]; we only vary the latent dimension $d$ and weight $\beta$ of the KL-term. By default, we set the token dimension to 16 (i.e., we predict 16 means and variances per token) and $\beta=5\cdot 10^{-5}$.
GIVT  For GIVT-Causal, we follow the original transformer decoder architecture [67] in decoder-only mode, but remove biases from attention layers, MLP blocks, and LayerNorms, and replace ReLU by GELU as is common practice. For GIVT-MaskGIT, we simply remove the attention mask during training and feed masked inputs instead of shifted ones. Unless otherwise stated, we use the BERT-Large configuration [12]. We configure our GIVT models to predict a single Gaussian distribution by default, but also explore predicting a mixture of 4 distributions per dimension, and predicting a multivariate Gaussian modeling the full covariance matrix of the tokens (instead of assuming a diagonal covariance matrix). For the conditional generation experiments, we use a learned embedding which we prepend to the embedded token sequence. To train GIVT, we use the AdaFactor optimizer from [73] with a cosine schedule (500 epochs; with linear warmup of 50 epochs), set the learning rate and weight decay to $10^{-3}$ and $10^{-4}$, respectively, the optimizer $\beta_{2}$ parameter to $0.95$, the dropout probability to $0.2$ for GIVT-causal and $0.4$ for GIVT-MaskGIT, and the batch size to $8192$. We use the same data augmentation as during VAE training (see [6, 43]), and sample from the VAE encoder distribution for every batch (an additional source of randomness besides data augmentation).
We implement GIVT in JAX [4] and use the JAX version of TensorFlow Probability [14] to implement the predicted distributions and compute the log-probabilities. We train GIVT-Casual and GIVT-MaskGIT models for $256\times 256$ ImageNet, and GIVT-MaskGIT additionally on $512\times 512$. We note that our VAE is trained on $256\times 256$ images, and we use it also for $512\times 512$ experiments (like [6]).
Figure 7: First two plots: Showing the impact of DB-CFG (Sec. 3.3.2) on sampling FID of our class-conditional $256\times 256$ models. First, we show our GIVT-Causal, with three sampling temperatures $t$, and see a compounding effect of $t$ and DB-CFG. Then, we show our GIVT-MaskGIT, using different inference schedules, and observe that apart from a linear schedule, all explored schedules achieve similar quality. Right-most plot: Average standard deviation predicted by GIVT-MaskGIT: when using CFG, we perform inference both with class labels (conditional) and without (unconditional). We plot the average predicted scale over 24 samples for class 94.
GIVT-MaskGIT inference  We fix the number of inference steps to 16 (as in [6]), and explore various inference schedules which define how many tokens are uncovered in each inference step. We employ the cosine schedule from the paper, as well as various power schedules (i.e. letting $r=i/16$ at step $i$, the fraction of masked tokens is given by $\cos(\pi/2r)$ for the cosine schedule, and by $1{-}r^{\alpha}$ for the pow:$\alpha$ schedule). As in [6], we sort tokens by likelihood at each step and sample using a “choice temperature” $t_{C}$.
Exploring the VAE latent space  To better understand the interplay between the feature dimension, the strength of the KL regularization $\beta$, the reconstruction quality of the VAE, and the sampling quality of the GIVT, we train VAEs with latent dimension in $\{4,8,16,32\}$ and $\beta$ in $\{2.5\cdot 10^{-5},5\cdot 10^{-5},10^{-4},2\cdot 10^{-4}\}$ using the VAE-training setup described at the beginning of this section. For each of the resulting VAEs we train a causal GIVT with the smaller BERT-Base [12] dimensions.
Evaluation  For the VAEs we report “reconstruction FID”, the FID obtained when reconstructing the 50k ImageNet validation images. For our GIVT variants and baselines, we report the sampling FID [21] when sampling a balanced set of 50k images covering all ImageNet classes. In both cases, we rely on the well-established ADM TensorFlow Suite [13], which uses the entire ImageNet training set as a reference. Furthermore, we also report Precision and Recall [52]. Finally, we evaluate the representation learning capabilities by training a linear classifier on an average-pooled intermediate representation of unconditional GIVT-Causal as in prior work [8, 71] (see App. D for details).
4.2 Panoptic segmentation and depth estimation
We build on the UViM framework [30], which uses a VQ-VAE to compress the label space of computer-vision dense-prediction tasks, and an encoder-decoder transformer taking the RGB image as an input and predicting the associated dense labels as discrete codes in the VQ-VAE latent space. Here, we replace the VQ-VAE with a $\beta$-VAE and use a GIVT encoder-decoder to model the continuous latent code. For the VAE, we use the same transformer-based autoencoder architecture (6-layer encoder and 12-layer decoder) and cross-entropy loss as [30]. We set $d=16$ and KL weight $\beta=2.5\cdot 10^{-4}$ for panoptic segmentation and $\beta=2\cdot 10^{-4}$ for depth estimation. To build an encoder-decoder GIVT model the same way as in [30], we employ the causal variant described for ImageNet generation and insert a cross-attention layer after each self-attention layer. Following [30], we use the ImageNet-21k-pretrained ViT-L/16 from [56] as the encoder, set the image resolution to 512px, and adopt the preprocessing and optimization hyper-parameters from [30]. We use the UViM variant without encoder and decoder context [30]. Finally, we consider variance scaling and beam search, selecting the best parameters on a held-out subset of the training set as in [30].
5 Results
5.1 Image generation
VAE latent space  In Fig. 4 we show how varying the weight $\beta$ of the KL term affects (i) the VAE reconstruction FID and (ii) the sampling FID of a Base-size GIVT-Causal trained on the corresponding latent sequence. For (i), increasing $\beta$ leads to worse reconstruction FID since the VAE can store less information in the latent. It shifts more of the modeling effort to the VAE decoder, so that the decoder becomes gradually more generative, which affects sampling quality (see [61, Sec., 7] for an extensive discussion).
For (ii), we see the opposite trend: increasing $\beta$ leads to decreased (better) sampling FID for GIVT models trained on the latent sequence. Arguably, this is because the VAE latent sequence more closely follows the Gaussian prior, and hence becomes easier for the GIVT to model. Note, however, that reconstruction FID lower-bounds the sampling FID, so one should chose a VAE with reconstruction FID smaller than the desired sampling FID. For this reason, we chose $\beta=5\cdot 10^{-5}$ for the VAE used in our main models, which obtains a reconstruction FID of 3.66 (note that our best GIVT-Causal model has sampling FID 4.29, see Table 1). Finally, we see that the latent dimension $d$ has quite a small impact on reconstruction and sampling FID.
Sampling FID  In Table 1 we show the sampling FID for three class-conditional model classes on $256\times 256$ ImageNet: GANs, diffusion-based approaches, and sequence modeling approaches. One can observe that our GIVT-Causal models outperform the quantization-based VQ-GAN in terms of FID, even though VQ-GAN has more than 4×4$\times$4 × the model parameters. We study the impact of our DB-CFG for both our models in Fig. 7. On the left, for causal models, we see the impact of temperature scaling, and obtain the best results with $t=0.95$. We did not observe benefits from temperature scaling for our GIVT-MaskGIT models, possibly due to the high-confidence selection-scheme already favoring high-certainty tokens. In the middle of Fig. 7, we instead show the impact of different inference schedules on GIVT-MaskGIT. We find that the cosine schedule used in the original VQ-VAE-based MaskGIT yields the best FID (by a small margin). Moving on to $512\times 512$ ImageNet, we show Sampling FID of available models in Table 2. We see that our GIVT-MaskGIT scales to this resolution, and obtains an FID of 5.81. Finally, we present unconditional results in App. C. This task is considerably harder, but GIVT outperforms the diffusion-based ADM [13] by a solid margin.
Samples  Fig. 2 shows 10 samples from GIVT-Causal, and App. E shows 10 samples for the GIVT-MaskGIT. We can see that the model produces high-fidelity, coherent samples. To study sample diversity, we sample multiple images for a fixed class in App. E.
In Fig. 6, one can see two samples from our VAE (obtained by decoding latents sampled from the prior), which show a soup of image textures. We then show different steps of the GIVT-MaskGIT inference, and observe similar behavior as in the VQ-based model ([6, Fig. 2]).
[…]
ADM [13] 250 $\numprint{23.2}$0
ADM-G [13] CG $=1.0$ 250 $\numprint{7.72}$
DiT-XL/2 [46] 250 $\numprint{12.03}$
DiT-XL/2-G [46] CFG $=1.5$ 250 $\numprint{3.04}$
[…]
Table 2: Results on class-conditional $512{\times}512$ ImageNet. We use the standard ADM evaluation suite for metrics, where FID is calculated w.r.t. the training set. Our masked GIVT-MaskGIT obtains competitive FID scores with only 16 inference steps. ${}^{\dagger}$umbers obtained by us from public code.
Ablations and visualizations  We did not observe clear benefits from using more than one mixture component per latent dimension in our image generation experiments (both for causal and MaskGIT-based GIVTs). This is perhaps unsurprising since our VAE’s latent codes are regularized towards a unimodal Gaussian prior. Using multiple mixtures might be more beneficial when, e.g., training GIVT on less-regularized latent spaces such as those of vision encoders pretrained in supervised or contrastive settings.
Further, predicting the full covariance matrix per latent vector, rather than assuming a diagonal covariance, only led to very modest gains of about 3% for MaskGIT-based GIVT. On the other hand, a full covariance matrix makes DB-CFG less tractable than the diagonal covariance (because the high dimensional multivariate distribution has more regions of low density).
In Fig. 7 (right), we visualize the predicted standard deviations as a function of the GIVT-MaskGIT inference step. The standard deviation gradually decreases, meaning that the predictions later in the sampling process become more certain. Furthermore, the unconditional predictions generally have a higher standard deviation, as expected.
Representation learning  Table 3 shows the ImageNet linear probing accuracy of unconditional GIVT-Causal and generative models from the literature (we chose the model variants closest in terms of model size and compute). GIVT-Causal matches VIM+ViT (ViT-VQ-GAN) [71] which has more than 2×2$\times$2 × the model parameters and 4×4$\times$4 × the sequence length (and hence FLOPs). GIVT-Causal is only outperformed by MAGE [36], whose latent encoder-decoder architecture is better suited for representation learning than decoder-only models. An investigation of the probing accuracy as function of the layer index can be found in App. D.
Model Type #Tok. #Param. Acc.$\uparrow$
BigBiGAN [15] 344M 61.3
iGPT-L [8] dec.-only 1024 1362M 60.3
VIM+CNN [71] dec.-only 1024 650M 61.8
VIM+ViT [71] dec.-only 1024 650M 65.1
MAGE ViT-L [36] enc.-dec. 256 404M 78.9
GIVT-Causal (Ours) dec.-only 256 304M 65.1
Table 3: ImageNet linear probing accuracy of unconditional GIVT-Causal and generative models from the literature. GIVT-Causal matches VIM+ViT (ViT-VQ-GAN) [71] which has more than 2×2$\times$2 × the model parameters and 4×4$\times$4 × the sequence length (and hence FLOPs). Type: (Latent) generative model type. #Param.: Number of parameters of the full (latent) generative model.
5.2 Panoptic segmentation and depth estimation
Table 4 compares the performance of a GIVT-based UViM variant with a VQ-VAE-based baseline (both without encoder/decoder context) on COCO Panoptic 2017 [29] and NYU Depth v2 [55]. We report the panoptic quality metric (PQ) [29] and RMSE, respectively, and find that our GIVT-based model outperforms the baseline by a small margin in panoptic segmentation and performs slightly worse in depth estimation. In App. E we show visual results.
[…]
recon. inference  recon. inference
UViM [30] 66.0 39.0 0.183 0.459
GIVT (ours) 71.0 40.2 0.195 0.474
Table 4: UViM based on GIVT-Causal and VQ-VAE evaluated on panoptic segmentation (COCO Panoptic 2017) and depth estimation (NYU Depth v2). We report the panoptic quality (PQ) and RMSE for the VAE/VQ-VAE reconstructions of the validation set label maps (recon.) and the inference metrics on the actual dense prediction tasks (inference). GIVT obtains metrics comparable to the VQ-based UViM.
6 Conclusion
In this paper, we proposed two small modifications to standard transformer decoder-only models to make them suitable to generating real-valued vectors. To our knowledge, this is the first decoder-only model amenable to generating sequences of real-valued vectors. In the context of image generation with VQ-GAN or Mask-GIT, this side-steps training difficulties such as low codebook usage in VQ-VAEs and corresponding mitigations like entropy losses or codebook-splitting algorithms, by enabling the use of standard VAEs which are much easier to train. Furthermore, our method avoids large embedding matrices because the feature representations can directly be consumed and predicted by our GIVT model. Surprisingly, our simple, quantization-free approach outperforms its VQ-based equivalent when using causal transformers for class-conditional image generation, and obtains strong performance in dense prediction tasks when applied to UViM. We hope that future work explores applications of GIVT to other modalities such as audio and time-series modeling.
Acknowledgments
We would like to thank André Susano Pinto, Neil Houlsby, Eirikur Agustsson, and Lucas Theis for inspiring discussions and helpful feedback on this project. We also thank Han Zhang for support with the VAE training code.
Appendix A Training details
A.1 Loss function and implementation details
We start with the case where GIVT models the distribution of each feature vector or soft token in the sequence independently across channels with a scalar $k$-mixture GMM, conditionally on all previous feature vectors in the sequence. The predicted GIVT output for a mini batch of size $B$ of $d$-dimensional feature sequences of length $L$ is a $B\times L\times 3dk$ tensor $y=[m;s;\pi]$ where $m,s$ and $\pi$ are $B\times L\times dk$ tensors stacked along the last dimension comprising means, standard deviations, and mixture weights, respectively. We assume the entries of $s$ to be lower-bounded by a small positive constant $\epsilon=10^{-5}$ (e.g., by applying a softplus function to the corresponding network outputs and clipping to $\epsilon$) and the entries of $\pi$ to be non-negative. $m$ is composed as $[m^{(1)};\ldots;m^{(k)}]$ where the $m^{(n)}$ are the $B\times L\times d$ mean tensors of the GMM components stacked along the last axis; $s$ and $\pi$ are composed in the same way. The mixture weights are assumed normalized across components (e.g., via a softmax function), i.e., $\sum_{n=1}^{k}\pi_{b,\ell,c}^{(n)}=1$ for all $b,\ell,c$.
Then, the negative log-likelihood of the feature sequences $\{z_{b}\}_{b=1}^{B}$ (i.e. the training images after embedding via the VAE encoder and applying reparametrization) w.r.t. the distribution $\tilde{p}$ predicted by GIVT can be written as
$-\sum_{b=1}^{B}\log(\tilde{p}(z_{b}))$
$=-\sum_{b=1}^{B}\log\left(\prod_{\ell=1}^{L}\prod_{c=1}^{d}\left(\sum_{n=1}^{k}\pi_{b,\ell,c}^{(n)}\mathcal{N}(z_{b,\ell,c}|m_{b,\ell,c}^{(n)},s_{b,\ell,c}^{(n)})\right)\right)$
$=-\sum_{b=1}^{B}\sum_{\ell=1}^{L}\sum_{c=1}^{d}\log\left(\sum_{n=1}^{k}\pi_{b,\ell,c}^{(n)}\mathcal{N}(z_{b,\ell,c}|m_{b,\ell,c}^{(n)},s_{b,\ell,c}^{(n)})\right),$ (2)
where $\mathcal{N}(z|m,s)$ is a Gaussian density
$\mathcal{N}(z|m,s)=\frac{1}{s\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{z-m}{s}\right)^{2}}.$
When using a single Gaussian instead of a GMM to model each output channel (i.e. $k=1$, as in most experiments) Eq. (2) becomes
$\sum_{b=1}^{B}\sum_{\ell=1}^{L}\sum_{c=1}^{d}\frac{1}{2}\left(\frac{z_{b,\ell,c}-m_{b,\ell,c}}{s_{b,\ell,c}}\right)^{2}+\log(s_{b,\ell,c})+\frac{1}{2}\log(2\pi).$
While this loss simplifies to a sum over the sequence dimension $\ell$, note that $m_{b,\ell,c},s_{b,\ell,c}$ (and in the GMM case $\pi_{b,\ell,c}$) are predicted from $z_{b,1},\ldots,z_{b,\ell-1}$ by GIVT (setting $z_{b,0}$ to a learned [CLS] or [BOS] vector). Further, it can be seen that lower-bounding the $s_{b,\ell,c}$ with $\epsilon>0$ avoids extremely large loss values when some $z_{b,\ell}$ fall into low-density regions of $\tilde{p}$.
For the multivariate case, where dependencies between feature channels are modeled, GIVT predicts a $d\times d$ covariance matrix $S^{(n)}_{b,\ell}$ for each mixture component $n$, and data negative log-likelihood becomes
$-\sum_{b=1}^{B}\sum_{\ell=1}^{L}\log\left(\sum_{n=1}^{k}\pi_{b,\ell}^{(n)}\tilde{\mathcal{N}}(z_{b,\ell}|m_{b,\ell}^{(n)},S_{b,\ell}^{(n)})\right),$
where $\tilde{\mathcal{N}}$ is a multivariate Gaussian density.
For all loss variants, the MaskGIT version of GIVT the sum over $L$ is reduced to the sum over the indices $l$ correspond to masked locations, and the loss is normalized by the number of mask tokens.
A.2 Image generation
For the image generation experiments on ImageNet, we adapt the CNN-based VQ-GAN tokenizer from MaskGIT (see vqgan_tokenizer.py). We replace the VQ layer and related losses with a Gaussian reparametrization layer [27], and we use the hyper parameters given in Table 5 for the CNN. For our transformers, we use the standard pre-LN setup (see, e.g., UViM vtt.py for a concrete implementation). We use the BERT-Large [12] configuration for the main models (24 layers, 16 attention heads, MLPs with hidden dimension 4096, embedding dimension 1024).
embedding_dim 32
filters 128
num_res_blocks 2
channel_multipliers (1, 1, 2, 2, 4)
conv_downsample False
activation_fn “swish”
norm_type “GN”
Table 5: Hyperparameters of the ImageNet CNN-based VAE encoder/decoder. Note that the 32 features of the embedding are split into 16 means and 16 scales, so our actual representation has $d=16$ channels after reparametrization.
ImageNet preprocessing  We preprocess the ImageNet data as follows, following prior work [6, 43]:
  • Decode JPEGs
  • Random crop such that between 80% and 100% of the source image are retained
  • Resize to target resolution ($256{\times}256$ or $512{\times}512$) using bicubic filters with anti-aliasing
  • Randomly flip the image horizontally with probability 0.5
A.3 Panoptic Segmentation and Depth Estimation
For our UViM panoptic segmentation and depth estimation experiments, we adopt the public UViM GitHub repo and only replace the VQ layer in stage I, adapt the transformer decoder in stage II, and modify the losses accordingly.
Appendix B DB-CFG implementation
We show the JAX implementation of the rejection sampler we use for DB-CFG in Fig. 9.
Appendix C Unconditional image generation
In Table 6 we present FID results for unconditional ImageNet generation.
[…]
ADM [13] $\numprint{26.2}$0
[…]
GIVT-Causal (Ours) $\numprint{24.95}$
[…]
Table 6: Results for unconditional $256\times 256$ ImageNet generation. We use the standard ADM evaluation suite for metrics, where FID is calculated w.r.t. the training set. We obtained samples for MAGE [36] using their GitHub code. Note that the MAGE Base and Large models have considerably more model parameters than our corresponding models because MAGE has a latent encoder-decoder model (rather than a decoder-only model).
Appendix D Probing intermediate representations
We extract intermediate representations form a GIVT-causal trained for $256\times 256$ unconditional ImageNet generation by average-pooling the output of a given layer over the sequence dimension (which results in a feature vector of dimension 1024 for L-sized models). Thereby, the input image is first encoded with the VAE encoder and then fed to GIVT as during teacher forcing (i.e. the latent sequence is first right-shifted and padded, and then fed to GIVT with causal attention mask).
Following prior work [8, 71], we investigate the impact of the layer index on the linear probing accuracy. To speed up the evaluation we rely on 25-shot classification with the fast evaluator form [73] and test a range of diverse data sets. The results presented in Fig. 8 show that the range of layers 10 to 20 lead to high accuracy for all data sets.
We choose layer 18 for the linear probing experiments using the full ImageNet training set. To select hyperparameters we set aside 1% of the training data and consider the selection of schedules and hyperparameters described in [62, App. A]. The results, along with other generative models from the literature, can be found in Table 3. See Sec. 5.1 for a discussion of these results.
Figure 8: 25-shot linear probing accuracy as a function of the GIVT-Causal layer trained on unconditional $256\times 256$ ImageNet. The range of layers 10 to 20 lead to high accuracy for all data sets.
Appendix E Additional visual examples
In Fig. 10, we show reconstructions from our VAE when feeding it images from the ImageNet validation set. We see low-level texture variations in the reconstructions when sampling from the encoder distribution multiple times. Fig. 11 shows multiple samples for a fixed label to show the diversity of our model and compare with baselines. In Fig. 12 we show various samples from MaskGIT for the same classes that we shown in Fig. 2. Fig. 13 shows samples from our $512\times 512$ GIVT-MaskGIT model. We explore changing labels midway through sampling (i.e., after generating the top half of the image) for GIVT-Causal in Fig. 14.
For our GIVT-Casual UViM models and VQ baselines, we show visual outputs in Figs. 15 16.
import jax
import jax.numpy as jnp
import tensorflow_probability as tfp
tfd = tfp.substrates.jax.distributions
# We assume the ps have shape (b, seq_len, c).
def rejection_sample(
    seed, p_cond: tfd.Normal, p_uncond: tfd.Normal, w, max_samples=1_000
  rng_sample, rng_uni = jax.random.split(seed, 2)
  scale_simple = jnp.stack([p_cond.scale, p_uncond.scale], -1).max(-1) * 2
  simple = tfd.Normal(loc=p_cond.loc, scale=scale_simple)  # Proposal.
  def unnormalized_pcfg(x):
    return jnp.exp((1 + w) * p_cond.log_prob(x) - w * p_uncond.log_prob(x))
  # Find scaling factor by checking for max around the conditional loc.
  points = p_cond.loc[None, ...] + jnp.linspace(-10, 10, 1001).reshape(
      -1, 1, 1, 1)
  fac = jnp.max(unnormalized_pcfg(points) / simple.prob(p_cond.loc), axis=0)
  # Shape (max_samples, b, seq_len, c)
  xs = simple.sample(seed=rng_sample, sample_shape=(max_samples,))
  facq = fac * simple.prob(xs)
  ys = jax.random.uniform(rng_uni, shape=facq.shape, minval=0.0, maxval=facq)
  p = unnormalized_pcfg(xs)
  # Mask is True if the sample is valid.
  mask = ys < p
  # Now we need to do tricks to get the first element in ‘mask‘ that is
  # True, in a jit-able way. We do this by making a shifted mask that is
  # False for every element after the first True. Example:
  # mask            [0, 1, 0, 1, 0, 0, 1, 0]
  # cmask           [0, 1, 1, 1, 1, 1, 1, 1]
  # shifted_cmask   [0, 0, 1, 1, 1, 1, 1, 1]
  # keep            [0, 1, 0, 0, 0, 0, 0, 0]  # <- picks the first valid!
  cmask = jnp.cumsum(mask, axis=0).astype(jnp.bool_)
  shifted_cmask = jnp.pad(
      cmask, [(1, 0), (0, 0), (0, 0), (0, 0)], constant_values=False
  )[:-1]
  assert shifted_cmask.shape == mask.shape
  keep = jnp.logical_and(cmask, jnp.logical_not(shifted_cmask))
  sample = jnp.where(keep, xs, 0).sum(0)  # Grab the first valid sample.
  ok = mask.sum(0) > 0  # Fall back to pdf_c if not ok.
  return jnp.where(ok, sample, pdf_c.sample(seed=rng_sample))
Figure 9: jax.jit-compatible implementation of the rejection sampler used for DB-CFG
Figure 10: Samples from the VAE encoder distribution: The first column is an image from the ImageNet validation set. We encode it with the VAE encoder, obtain the approximate posterior distribution, and sample from it 4 times. The resulting 4 latent sequences are then decoded to images (last 4 columns). The semantic layout of the input image is well preserved, only the low-level textures change.
GIVT-Causal (Ours)
GIVT-MaskGIT (Ours)
Validation Set
BigGAN [5] (via [6, Fig. 9])
VQ-based MaskGIT [6, Fig. 9]
Figure 11: Various samples for the 980 class to show the diversity of our samples. For reference, we copy examples from from [6, Fig. 9].
Figure 12: Selected $256\times 256$ samples from GIVT-MaskGIT, for the same 10 ImageNet classes as in Fig. 2.
Figure 13: Selected $512\times 512$ samples from the GIVT-MaskGIT, for the same 10 ImageNet classes as in Fig. 2.
Figure 14: Changing the label half way through sampling (i.e., after the top half is generated) for GIVT-Causal: The first column uses the same label for the top and bottom half, the other columns switch to a different labels. Top row labels: “golden retriever” (207) for the top half; “otter” (360), “gorilla” (366), “camel” (355) for the bottom half. Bottom row labels: “bird house” (448) for the top half; “boat house” (449), “light house” (437), “bakery” (415) for the bottom half. Note that for each row, the top row latent is always the same, but the overall color balance in the RGB output might be different because of the VAE decoder (possibly due to GroupNorm layers).
Input
Ground Truth
GIVT-Causal UViM
VQ-based UViM
Figure 15: ‘COCO Panoptic Segmentation’ visual examples from our UViM models. Panoptic segmentation visual examples from COCO Panoptic 2017 for UViM based on GIVT-Causal and standard VQ-based UViM.
Input
Ground Truth
GIVT-Causal UViM
VQ-based UViM
Figure 16: Depth estimation visual examples from NYU Depth v2 for UViM based on GIVT-Causal and standard VQ-based UViM.